# FSFSplatter: Build Surface and Novel Views with Sparse-Views within 3min

### v1

@@ Abstract
Gaussian Splatting has become a leading reconstruction technique, known for its high-quality novel view synthesis and detailed reconstruction. However, most existing methods require dense, calibrated views. Reconstructing from free sparse images often leads to poor surface due to limited overlap and overfitting. We introduce FSFSplatter, a new approach for fast surface reconstruction from free sparse images. Our method integrates end-to-end dense Gaussian initialization, camera parameter estimation, and geometry-enhanced scene optimization. Specifically, FSFSplatter employs a large Transformer to encode multi-view images and generates a dense and geometrically consistent Gaussian scene initialization via a self-splitting Gaussian head. It eliminates local floaters through contribution-based pruning and mitigates overfitting to limited views by leveraging depth and multi-view feature supervision with differentiable camera parameters during rapid optimization. FSFSplatter outperforms current state-of-the-art methods on widely used DTU and Replica.
@@ 1 Introduction
    We propose FSFSplatter, a fast surface reconstruction algorithm from free-sparse images. Compared to prior works, our method achieves accurate reconstruction results both with and without per-scene optimization.
@@
    We demonstrate state-of-the-art (SOTA) performance on widely-used DTU and Replica datasets upon surface reconstruction and NVS.
@@ 3 Method
We propose a fast reconstruction method based on free-sparse RGB images that enables robust 3D surface reconstruction from only three input images within 3 minutes.
@@ 4.1 Settings
We employ widely adopted metrics including PSNR, SSIM[38], and LPIPS[39] to evaluate NVS quality, while utilizing Chamfer Distance (CD)[35] to evaluate surface reconstruction quality.
@@
We evaluate our method on DTU datasets [35] and Replica datasets [37]. DTU datasets have been extensively used in prior surface reconstruction works. We evaluate on all 15 scenes from DTU, where each scene contains 49 or 64 images at a resolution of $1600\times 1200$. Following previous works [23, 41], we select views 23, 24, and 33 as input and downsample the images to $800\times 600$. Since DTU datasets primarily consist of object-level scenes, we further validate our algorithm on Replica datasets to assess its generalization capability and performance on scene-level scenes. We select the first five scenes and randomly choose three input views at a resolution of $1200\times 680$.
@@ 4.2 Evaluation of Surface Reconstruction
We first evaluate the surface reconstruction accuracy (CD) under different settings across all baselines. As shown in Table1, Table2, and Figure5, our method demonstrates significant advantages in CD error even when compared with per-scene optimization methods using their default iteration numbers, and remains competitive with methods using ground truth camera parameters as input. Particularly in scene-level scenarios, our method exhibits superior robustness. The error is reduced by at least 28.39% and 34.37% on DTU and Replica datasets respectively. Even without per-scene optimization, the surface reconstruction error is reduced by at least 14.62% and 21.93% respectively.
@@
Table 1: Quantitative experiments of CD on DTU datasets under different settings. We conduct experiments on all 15 scenes.
@@
Table 2: Quantitative experiments of CD on Replica datasets under different settings. We conduct experiments on 5 scenes.
@@ 4.3 Evaluation of NVS
Subsequently, we evaluate the performance of our method on NVS. As demonstrated in Table 1 and Table 2, under conditions of sparse viewpoints and ground-truth camera parameters, simply increasing the number of iterations does not yield better reconstruction results; conversely, it leads to catastrophic scene geometry due to overfitting on the limited views. Therefore, we evaluate the NVS performance under equal iteration numbers. As shown in Table 3, the NVS performance of our method significantly surpasses other approaches. Even without per-scene optimization, where the scene is generated solely through end-to-end dense Gaussian generation, our method still demonstrates a clear advantage in NVS performance. Moreover, compared to the overfitting on limited views observed in methods represented by 3DGS, our approach effectively preserves scene geometry while performing per-scene optimization. The NVS error (LPIPS as an example) is reduced by at least 46.19% on DTU datasets and 73.13% on Replica datasets. Even without per-scene optimization, the error is decreased by at least 37.14% and 1.29% respectively.
@@
Table 3: Quantitative evaluation of NVS on DTU and Replica datasets under different settings.
@@ 4.4 Ablation Study
To validate the effectiveness of each component in our proposed algorithm, we conducted comprehensive ablation studies. Specifically, we performed ablations on both the end-to-end dense Gaussian initialization and the geometrically-enhanced scene optimization, with the surface reconstruction quality evaluated using CD. Experiments were carried out on the DTU and Replica datasets. For per-scene optimization, we ablated the following components: monocular depth supervision $\mathcal{L}_{\text{rank}}$, multi-view feature consistency supervision $\mathcal{L}_{\text{MVS}}$, and differentiable camera poses $T^{\text{cam}}_{k}$. For Gaussian initialization, we ablated the contribution-based Gaussian pruning $Trim$ and the self-splitting-based Gaussian scene densification $\mathbb{D}$.
@@ 5 Conclusion
We propose FSFSplatter, a novel algorithm for rapid novel view synthesis and surface reconstruction from free-sparse RGB images. Our approach integrates Gaussian scene densification initialization and camera parameter estimation into an end-to-end large Transformer, which employs a self-splitting Gaussian head to map semi-dense point cloud features into a dense Gaussian scene, while applying contribution-based Gaussian pruning to eliminate useless floating artifacts. To address the issue of local view overfitting caused by sparse views, we incorporate supervision from monocular depth estimation and multi-view feature consistency to regularize the scene. Furthermore, we parameterize camera poses as differentiable variables to mitigate errors from camera parameter estimation. Compared to prior methods, FSFSplatter demonstrates superior accuracy and robustness, both with and without per-scene optimization. We achieve SOTA performance in free-sparse view surface reconstruction and NVS for both object-level and scene-level scenes, validated on the widely used DTU and Replica datasets.

### v2

# FSFSplatter: Build Surface and Novel Views with Sparse-Views within 2min

@@ Abstract
Gaussian Splatting has become a leading reconstruction technique, known for its high-quality novel view synthesis and detailed reconstruction. However, most existing methods require dense, calibrated views. Reconstructing from free sparse images often leads to poor surface due to limited overlap and overfitting. We introduce FSFSplatter, a new approach for fast surface reconstruction from free sparse images. Our method integrates end-to-end dense Gaussian initialization, camera parameter estimation, and geometry-enhanced scene optimization. Specifically, FSFSplatter employs a large Transformer to encode multi-view images and generates a dense and geometrically consistent Gaussian scene initialization via a self-splitting Gaussian head. It eliminates local floaters through contribution-based pruning and mitigates overfitting to limited views by leveraging depth and multi-view feature supervision with differentiable camera parameters during rapid optimization. FSFSplatter outperforms current state-of-the-art methods on widely used DTU, Replica, and BlendedMVS datasets.
Figure 1: We address the inherent ambiguity and local view overfitting in the reconstruction of free-sparse images, proposing a novel pipeline named FSFSplatter that enables surface reconstruction and novel view synthesis within 2 minutes. FSFSplatter significantly reduces the number of iterations required while decreasing surface reconstruction error by at least 28.39% and novel view synthesis error by at least 46.19%. The figure illustrates the scene "scan63" from DTU datasets.
@@ 1 Introduction
    We propose FSFSplatter, a fast surface reconstruction algorithm from free-sparse images. Compared to prior works, FSFSplatter achieves accurate reconstruction results both with and without per-scene optimization.
@@
    We demonstrate state-of-the-art (SOTA) performance on widely-used DTU, Replica, and BlendedMVS datasets upon surface reconstruction and NVS.
@@ 3 Method
We propose a fast reconstruction method based on free-sparse RGB images that enables robust 3D surface reconstruction from only three input images within 2 minutes.
@@
Figure 2: Overview of FSFSplatter. We employ a large alternating attention network as the backbone and generate high-dimensional semi-dense Gaussian scenes through independent heads. These are subsequently mapped into a dense Gaussian scene via a self-splitting densify MLP, forming an end-to-end framework for dense Gaussian initialization and camera parameter estimation. Through differentiable Gaussian scene construction from initialization and joint supervision of depth, multi-view stereo features, and RGBs during the optimization process, scene optimization achieves geometric enhancement, effectively mitigating error surfaces caused by overfitting to free-sparse views.
@@ 4.1 Settings
We employ widely adopted metrics including PSNR, SSIM[38], and LPIPS[39] to evaluate NVS quality; employ Chamfer Distance (CD)[35] to evaluate surface reconstruction quality; employ ATE RMSE[40] (rotation and translation) to evaluate camera pose estimation.
@@
We evaluate FSFSplatter on DTU datasets [35], Replica datasets [37] and BlendedMVS datasets[34]. DTU datasets have been extensively used in prior surface reconstruction works. We evaluate on all 15 scenes from DTU, where each scene contains 49 or 64 images at a resolution of $1600\times 1200$. Following previous works [23, 42], we select views 23, 24, and 33 as input and downsample the images to $800\times 600$. Since DTU datasets primarily consist of object-level scenes, we further validate our algorithm on Replica and BlendedMVS datasets to assess its generalization capability and performance on scene-level scenes, which respectively contain simulated images and real-captured images. We respectively select the five scenes and randomly choose three input views at a resolution of $1200\times 680$ and $768\times 576$.
Experimental Platform
Experiments were carried out utilizing an NVIDIA RTX 5090 GPU alongside an Intel Platinum 8470Q CPU. The software setup operated on Ubuntu 20.04, with PyTorch version 2.3.1 and CUDA version 12.1.
@@ 4.2 Evaluation of Surface Reconstruction
We first evaluate the surface reconstruction accuracy (CD) under different settings across all baselines. As shown in Table1, Table2, and Figure7, FSFSplatter demonstrates significant advantages in CD error even when compared with per-scene optimization methods using their default iteration numbers, and remains competitive with methods using ground truth camera parameters as input. Particularly in scene-level scenarios, FSFSplatter exhibits superior robustness. The error is reduced by at least 28.39% and 34.37% on DTU and Replica datasets respectively. Even without per-scene optimization, the surface reconstruction error is reduced by at least 14.62% and 21.93% respectively.
@@
Table 1: Quantitative experiments of CD ($m$) on DTU datasets under different settings. We conduct experiments on all 15 scenes.
@@
Table 2: Quantitative experiments of CD ($m$) on Replica datasets under different settings. We conduct experiments on 5 scenes.
@@ 4.3 Evaluation of NVS
Subsequently, we evaluate the performance of FSFSplatter on NVS. As demonstrated in Table 1 and Table 2, under conditions of sparse viewpoints and ground-truth camera parameters, simply increasing the number of iterations does not yield better reconstruction results; conversely, it leads to catastrophic scene geometry due to overfitting on the limited views. Therefore, we evaluate the NVS performance under equal iteration numbers. As shown in Table 3 and Figure 8, the NVS performance of FSFSplatter significantly surpasses other approaches. Even without per-scene optimization, where the scene is generated solely through end-to-end dense Gaussian generation, FSFSplatter still demonstrates a clear advantage in NVS performance. Moreover, compared to the overfitting on limited views observed in methods represented by 3DGS, our approach effectively preserves scene geometry while performing per-scene optimization. The NVS error (LPIPS as an example) is reduced by at least 46.19% on DTU datasets, 73.13% on Replica datasets, and 87.35% on BlendedMVS datasets. Even without per-scene optimization, the error is decreased by at least 37.14%, 1.29%, and 14.36% respectively.
Figure 8: Visualization of NVS on DTU, Replica, and BlendedMVS datasets.
@@
[…]
Table 3: Quantitative evaluation of NVS on DTU, Replica, and BlendedMVS datasets under different settings.
4.4 Evaluation of pose extimation
As a surface reconstruction and NVS method based on free-sparse images, we also evaluated the accuracy of camera pose estimation on the Replica and DTU datasets. Since both FSFSplatter and the pipelines are based on RGB images, we aligned the scale by registering the camera poses following established works [4].
As shown in Table 4, FSFSplatter achieves competitive performance in terms of ATE RMSE compared to SOTA pose estimation methods, while simultaneously delivering high-quality surface reconstruction and NVS results.
CF-3DGS FreeSplatter Fast3R Dust3r VGGT Ours
[…]
Table 4: Pose estimation evaluation ATE RMSE on Replica and DTU datasets.
4.5 Evaluation of optimization speed
As a reconstruction method that supports per-scene optimization, FSFSplatter achieves high reconstruction accuracy from the initial stage due to its well-structured Gaussian dense initialization. Consequently, FSFSplatter attains remarkably short optimization iteration times. As shown in Table 5, we evaluate the optimization speed of FSFSplatter and other per-scene optimization-based reconstruction algorithms on Replica datasets.
Method Time(s)
3DGS 429.45
2DGS 840.37
FatesGS 915.93
PGSR 1018.2
Ours 107.39
Table 5: Optimization speed evaluation of per-scene optimization-based algorithms on Replica datasets.
@@ 4.6 Ablation study
To validate the effectiveness of each component in our proposed algorithm, we conducted comprehensive ablation studies. Specifically, we performed ablations on both the end-to-end dense Gaussian initialization and the geometrically-enhanced scene optimization, with the surface reconstruction quality evaluated using CD. Experiments were carried out on the DTU and Replica datasets. For per-scene optimization, we ablated the following components: monocular depth supervision $\mathcal{L}_{\text{rank}}$, depth smoothness loss $\mathcal{L}_{\text{smooth}}$, multi-view feature consistency supervision $\mathcal{L}_{\text{MVS}}$, and differentiable camera poses $T^{\text{cam}}_{k}$. For Gaussian initialization, we ablated the contribution-based Gaussian pruning $Trim$ and the self-splitting-based Gaussian scene densification $\mathbb{D}$.
@@ 5 Conclusion
We propose FSFSplatter, a novel algorithm for rapid novel view synthesis and surface reconstruction from free-sparse RGB images. Our approach integrates Gaussian scene densification initialization and camera parameter estimation into an end-to-end large Transformer, which employs a self-splitting Gaussian head to map semi-dense point cloud features into a dense Gaussian scene, while applying contribution-based Gaussian pruning to eliminate useless floating artifacts. To address the issue of local view overfitting caused by sparse views, we incorporate supervision from monocular depth estimation and multi-view feature consistency to regularize the scene. Furthermore, we parameterize camera poses as differentiable variables to mitigate errors from camera parameter estimation. Compared to prior methods, FSFSplatter demonstrates superior accuracy and robustness, both with and without per-scene optimization. We achieve SOTA performance in free-sparse view surface reconstruction and NVS for both object-level and scene-level scenes, validated on the widely used DTU, Replica, and BlendedMVS datasets.
Appendix A Details of framework and implementation
A.1 Implementation
The scene representation of FSFSplatter is based on 2DGS, and the rasterization pipeline is implemented based on AbsGS [43] and TrimGS [44]. Compared to basic 2DGS pipeline [2], the rasterization pipeline of FSFSplatter employs Abs accumulated gradient $\bar{\Delta}_{i}^{\text{ACC}}$ when caculating the accumulated gradient ${\Delta}_{i}^{\text{ACC}}$ (i.e., the gradient of the loss function $\mathcal{L}$ with respect to the projected pixel coordinate $u$). This approach mitigates gradient collision, thereby enabling accurate identification of suboptimal Gaussian primitives and addressing the blurring and loss of detail caused by over-reconstruction during Gaussian splitting, as shown in Equation S1. Here, $m$ denotes the sequence of all affected pixels. Contribution-based pruning is implemented following TrimGS, with the key distinction being the extension from pruning designed for NVS in 3DGS to pruning designed for surface reconstruction in 2DGS, as formulated in Equation 6.
$\bar{\Delta}_{i}^{\text{ACC}}=\sum_{j=1}^{m}\left|\frac{\partial\mathcal{L}_{j}}{\partial u_{i}}\right|,\quad{\Delta}_{i}^{\text{ACC}}=\sum_{j=1}^{m}\frac{\partial\mathcal{L}_{j}}{\partial u_{i}}$ (S1)
A.2 Parameter Setting
For the contribution-based Gaussian pruning, we perform a pruning at the start of optimization and subsequently repeat it every 500 iterations, removing the 5% of Gaussian primitives with the lowest contribution. For the loss function, the parameters $m$, $n_{1}$, and $n_{2}$ are set to $\times 10^{-4}, 1×1$$\times 10^{-2}$, and $1\times 10^{-4}$, respectively.
A.3 Evaluation Metrics
We employ PSNR, SSIM, and LPIPS to comprehensively evaluate the quality of NVS. For DTU datasets [35] and BlendedMVS datasets [34], where transparent masks are provided, we apply masks to all images prior to evaluation to ensure fair comparisons. To assess surface reconstruction quality, we utilize the Chamfer Distance (CD). Following previous works [35, 23], we adopt the L2 Chamfer Distance as our evaluation metric, as shown in Equation S2.
For evaluating camera pose estimation quality, we use the ATE RMSE. Since FSFSplatter is entirely based on RGB images, estimated poses differ from the ground truth by a scale factor. Following prior research [40], we perform rigid registration on the camera poses before evaluation and compute the absolute error in the ground truth scale. This approach eliminates inconsistencies in error evaluation caused by scale differences.
$\mathrm{CD}\left(\mathcal{M},\mathcal{M}_{\text{gt}}\right)=\frac{1}{2\left|\mathcal{M}\right|}\sum_{{x}\in\mathcal{M}}\min_{{y}\in\mathcal{M}_{\text{gt}}}\|{x}-{y}\|_{2}+\frac{1}{2\left|\mathcal{M}_{\text{gt}}\right|}\sum_{{y}\in\mathcal{M}_{\text{gt}}}\min_{{x}\in\mathcal{M}}\|{x}-{y}\|_{2}$ (S2)
Appendix B Detailed experiments
B.1 Detailed NVS quantization experiments
Table 3 presents the average quantitative evaluation results of NVS on DTU, Replica, and BlendedMVS datasets, while Tables S1, Table S2, and Table S3 provide the detailed per-scene evaluation results for each dataset.
Scan 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
PSNR 31.128 29.295 25.873 36.033 34.249 32.639 32.441 37.970 31.912 36.649 37.903 34.387 34.111 40.040 40.704 34.356
3DGS-GT SSIM 0.841 0.887 0.865 0.970 0.949 0.964 0.935 0.963 0.937 0.933 0.951 0.945 0.931 0.959 0.969 0.933
LPIPS 0.279 0.158 0.231 0.080 0.118 0.094 0.199 0.103 0.152 0.157 0.158 0.164 0.173 0.132 0.098 0.153
PSNR 20.265 22.157 23.860 28.613 26.711 25.606 26.693 30.251 25.402 28.982 31.817 27.723 30.078 36.417 35.506 28.006
2DGS-GT SSIM 0.609 0.782 0.745 0.907 0.927 0.937 0.882 0.941 0.869 0.895 0.923 0.903 0.886 0.940 0.953 0.873
LPIPS 0.489 0.230 0.326 0.154 0.144 0.134 0.276 0.130 0.232 0.204 0.193 0.216 0.226 0.158 0.118 0.215
PSNR 27.518 26.766 25.185 35.262 28.352 27.879 29.146 30.967 27.968 30.423 30.004 32.456 34.668 40.089 38.549 31.016
PGSR-GT SSIM 0.841 0.905 0.906 0.978 0.944 0.963 0.911 0.950 0.922 0.937 0.954 0.937 0.957 0.973 0.976 0.937
LPIPS 0.251 0.122 0.156 0.046 0.089 0.071 0.149 0.096 0.115 0.119 0.113 0.143 0.081 0.064 0.054 0.111
PSNR 25.790 17.563 14.710 26.701 14.782 24.567 23.939 17.852 20.589 17.613 27.563 25.467 27.786 30.111 29.991 23.002
FatesGS-GT SSIM 0.780 0.703 0.800 0.795 0.804 0.816 0.804 0.725 0.800 0.801 0.849 0.792 0.844 0.861 0.842 0.801
LPIPS 0.343 0.281 0.280 0.315 0.240 0.374 0.333 0.438 0.323 0.375 0.351 0.400 0.296 0.332 0.358 0.336
PSNR 20.938 23.568 25.674 28.973 29.863 28.280 27.256 34.471 26.401 32.361 30.170 28.248 30.539 36.792 37.067 29.373
3DGS SSIM 0.618 0.787 0.730 0.906 0.933 0.943 0.894 0.954 0.878 0.901 0.928 0.909 0.889 0.941 0.955 0.878
LPIPS 0.487 0.226 0.329 0.149 0.140 0.128 0.261 0.118 0.225 0.201 0.186 0.201 0.225 0.158 0.118 0.210
PSNR 20.212 16.725 21.856 28.686 26.678 25.487 26.283 27.593 22.665 28.945 30.618 26.986 30.025 30.796 35.225 26.585
2DGS SSIM 0.603 0.724 0.693 0.907 0.926 0.932 0.876 0.929 0.820 0.894 0.924 0.900 0.885 0.910 0.950 0.858
LPIPS 0.495 0.267 0.355 0.154 0.145 0.140 0.290 0.143 0.292 0.205 0.192 0.219 0.228 0.197 0.124 0.230
PSNR 21.146 17.712 21.858 29.532 26.743 24.684 26.983 27.803 22.494 28.793 30.880 28.184 31.726 32.199 36.279 27.134
PGSR SSIM 0.636 0.716 0.733 0.941 0.929 0.934 0.876 0.926 0.811 0.912 0.944 0.908 0.928 0.922 0.965 0.872
LPIPS 0.463 0.289 0.326 0.094 0.106 0.112 0.195 0.138 0.256 0.158 0.125 0.184 0.121 0.144 0.069 0.185
PSNR 16.564 16.781 20.542 23.955 29.153 20.238 23.836 24.370 20.296 22.916 25.261 24.491 26.948 31.785 29.240 23.758
FatesGS SSIM 0.506 0.729 0.662 0.828 0.947 0.889 0.798 0.894 0.783 0.833 0.876 0.867 0.844 0.911 0.916 0.819
LPIPS 0.592 0.273 0.376 0.210 0.090 0.159 0.323 0.170 0.287 0.244 0.231 0.244 0.238 0.168 0.141 0.250
PSNR 19.495 17.156 21.760 24.198 24.421 27.438 27.590 29.752 23.873 26.474 28.559 28.852 27.106 30.780 32.207 25.977
CF-3DGS SSIM 0.650 0.754 0.715 0.839 0.918 0.933 0.868 0.949 0.845 0.883 0.884 0.917 0.859 0.904 0.927 0.856
LPIPS 0.328 0.217 0.274 0.147 0.118 0.098 0.190 0.081 0.198 0.147 0.160 0.127 0.186 0.133 0.095 0.167
PSNR 15.427 8.386 16.189 15.365 22.240 7.930 21.433 25.158 14.965 21.129 24.063 19.481 10.203 25.891 26.146 18.267
FreeSplatter SSIM 0.381 0.566 0.399 0.577 0.822 0.658 0.652 0.851 0.586 0.732 0.739 0.718 0.548 0.776 0.827 0.655
LPIPS 0.434 0.304 0.379 0.238 0.139 0.246 0.233 0.111 0.311 0.228 0.207 0.211 0.322 0.169 0.132 0.244
PSNR 13.207 15.304 15.559 17.032 19.373 13.684 20.432 23.233 17.266 19.399 20.146 21.977 13.602 22.305 22.993 18.368
VGGT SSIM 0.511 0.735 0.619 0.744 0.868 0.827 0.744 0.898 0.754 0.807 0.803 0.836 0.723 0.832 0.860 0.771
LPIPS 0.404 0.204 0.282 0.170 0.130 0.173 0.206 0.100 0.214 0.176 0.194 0.164 0.238 0.146 0.107 0.194
PSNR 24.615 22.975 26.135 26.706 28.720 29.818 27.434 33.895 26.526 25.759 31.803 27.280 28.393 34.096 35.101 28.617
Ours(wo Opt.) SSIM 0.751 0.822 0.763 0.874 0.939 0.955 0.887 0.964 0.879 0.913 0.922 0.917 0.884 0.932 0.951 0.890
LPIPS 0.312 0.153 0.201 0.119 0.082 0.065 0.141 0.060 0.153 0.128 0.143 0.122 0.144 0.091 0.068 0.132
PSNR 25.909 25.376 27.419 26.491 33.835 32.678 31.306 28.887 28.919 27.382 33.239 32.483 30.114 34.646 35.083 30.251
Ours SSIM 0.802 0.860 0.800 0.872 0.988 0.961 0.929 0.945 0.910 0.899 0.925 0.952 0.902 0.937 0.949 0.909
LPIPS 0.231 0.133 0.173 0.104 0.015 0.060 0.100 0.090 0.121 0.149 0.117 0.080 0.124 0.082 0.067 0.110
Table S1: Detailed NVS experiments on DTU datasets.
Scan Office0 Office1 Office2 Room0 Room1 Mean
PSNR 24.503 29.560 18.541 18.452 20.942 22.400
3DGS-GT SSIM 0.848 0.883 0.749 0.786 0.755 0.804
LPIPS 0.417 0.348 0.456 0.428 0.393 0.409
PSNR 24.769 27.655 17.313 18.677 18.372 21.357
2DGS-GT SSIM 0.852 0.877 0.740 0.788 0.742 0.800
LPIPS 0.416 0.353 0.469 0.425 0.417 0.416
PSNR 24.867 28.187 18.165 18.630 20.535 22.077
PGSR-GT SSIM 0.793 0.880 0.701 0.739 0.753 0.773
LPIPS 0.450 0.293 0.494 0.457 0.398 0.418
PSNR 25.736 28.701 17.818 18.287 20.870 22.282
FatesGS-GT SSIM 0.863 0.875 0.747 0.783 0.761 0.806
LPIPS 0.400 0.337 0.461 0.426 0.388 0.402
PSNR 25.428 27.012 17.730 18.772 21.346 22.058
3DGS SSIM 0.861 0.861 0.745 0.785 0.758 0.802
LPIPS 0.411 0.357 0.464 0.434 0.391 0.411
PSNR 23.905 26.530 17.560 18.530 19.844 21.274
2DGS SSIM 0.850 0.871 0.742 0.783 0.749 0.799
LPIPS 0.415 0.354 0.462 0.433 0.398 0.412
PSNR 24.280 25.815 17.366 18.774 20.973 21.441
PGSR SSIM 0.780 0.871 0.691 0.743 0.750 0.767
LPIPS 0.465 0.301 0.509 0.463 0.404 0.428
PSNR 24.723 30.608 17.788 19.013 20.544 22.535
FatesGS SSIM 0.856 0.900 0.744 0.789 0.748 0.808
LPIPS 0.408 0.325 0.461 0.422 0.395 0.402
PSNR 22.460 24.403 12.632 15.656 16.582 18.347
CF-3DGS SSIM 0.778 0.734 0.603 0.726 0.636 0.695
LPIPS 0.336 0.314 0.477 0.427 0.382 0.387
PSNR 20.143 20.134 8.848 12.575 11.074 14.555
FreeSplatter SSIM 0.742 0.635 0.514 0.694 0.556 0.628
LPIPS 0.422 0.471 0.533 0.409 0.454 0.458
PSNR 16.141 20.113 10.272 9.489 9.219 13.047
VGGT SSIM 0.620 0.630 0.494 0.476 0.436 0.531
LPIPS 0.463 0.385 0.538 0.577 0.606 0.514
PSNR 24.230 28.125 19.985 22.502 22.711 23.511
Ours(wo Opt.) SSIM 0.779 0.773 0.720 0.773 0.743 0.758
LPIPS 0.403 0.368 0.419 0.373 0.349 0.382
PSNR 40.059 41.109 24.113 38.218 35.461 35.792
Ours SSIM 0.978 0.968 0.931 0.974 0.976 0.965
LPIPS 0.073 0.137 0.131 0.079 0.057 0.095
Table S2: Detailed NVS experiments on Replica datasets.
Scan 5a48d4 5a618c 5a69c4 5b908d 59338e Mean
PSNR 18.990 16.663 17.389 18.280 17.341 17.733
3DGS-GT SSIM 0.696 0.469 0.471 0.518 0.582 0.547
LPIPS 0.411 0.516 0.516 0.572 0.480 0.499
PSNR 17.751 17.420 16.283 17.722 17.027 17.241
2DGS-GT SSIM 0.687 0.469 0.462 0.520 0.576 0.543
LPIPS 0.418 0.512 0.522 0.570 0.483 0.501
PSNR 18.755 17.003 16.273 19.026 17.098 17.631
PGSR-GT SSIM 0.672 0.461 0.483 0.536 0.567 0.544
LPIPS 0.382 0.529 0.508 0.545 0.486 0.490
PSNR 16.678 16.324 16.423 19.991 13.396 16.563
FatesGS-GT SSIM 0.678 0.472 0.459 0.524 0.542 0.535
LPIPS 0.416 0.511 0.523 0.558 0.521 0.506
PSNR 24.075 16.587 18.005 18.310 16.788 18.753
3DGS SSIM 0.787 0.465 0.476 0.519 0.578 0.565
LPIPS 0.326 0.517 0.510 0.574 0.486 0.483
PSNR 20.773 17.093 16.617 18.195 16.371 17.810
2DGS SSIM 0.722 0.464 0.467 0.520 0.573 0.549
LPIPS 0.386 0.513 0.519 0.569 0.489 0.495
PSNR 22.861 16.725 17.730 19.215 16.742 18.655
PGSR SSIM 0.756 0.461 0.501 0.540 0.558 0.563
LPIPS 0.309 0.528 0.497 0.548 0.492 0.475
PSNR 20.626 14.752 18.949 19.419 17.816 18.312
FatesGS SSIM 0.772 0.418 0.498 0.526 0.612 0.565
LPIPS 0.333 0.537 0.496 0.552 0.466 0.477
PSNR 9.222 14.539 12.652 12.736 15.536 12.937
CF-3DGS SSIM 0.509 0.387 0.398 0.334 0.533 0.432
LPIPS 0.451 0.454 0.465 0.517 0.378 0.453
PSNR 16.111 16.398 16.396 14.015 13.026 15.189
FreeSplatter SSIM 0.629 0.415 0.451 0.493 0.452 0.488
LPIPS 0.390 0.503 0.508 0.521 0.533 0.491
PSNR 11.565 15.799 12.832 14.345 12.493 13.407
VGGT SSIM 0.590 0.543 0.478 0.478 0.546 0.527
LPIPS 0.403 0.347 0.425 0.457 0.425 0.411
PSNR 17.721 21.681 17.158 21.751 19.262 19.515
Ours(wo Opt.) SSIM 0.700 0.670 0.551 0.604 0.644 0.634
LPIPS 0.325 0.268 0.383 0.424 0.362 0.353
PSNR 28.986 32.801 29.437 33.096 29.362 30.737
Ours SSIM 0.977 0.973 0.951 0.964 0.957 0.965
LPIPS 0.038 0.034 0.065 0.055 0.068 0.052
Table S3: Detailed NVS experiments on BlendedMVS datasets.
B.2 Detailed optimization speed experiments
Table 5 presents the average time consumption of optimization-based reconstruction algorithms on Replica datasets, while Tables S4 provide the detailed per-scene evaluation results.
3DGS 2DGS FatesGS PGSR FSFSplatter
Office0 407.72 800.88 638.38 873.45 105.19
Office1 383.28 938.22 975.06 1102.7 103.22
Office2 438.57 825.10 811.89 1037.2 112.44
Room0 434.96 775.38 838.49 853.54 106.88
Room1 482.75 862.31 1315.8 1224.1 109.23
Average 429.45 840.37 915.93 1018.1 107.39
Table S4: Detailed optimization speed experiments on Replica datasets.
B.3 Detailed pose estimation experiments
Table 4 presents the average ATE RMSE error on DTU and Replica datasets, while Tables S5 provide the detailed per-scene evaluation results.
CF-3DGS FreeSplatter Fast3R Dust3r VGGT Ours
Rot Trans Rot Trans Rot Trans Rot Trans Rot Trans Rot Trans
Replica Office0 74.930 38.575 51.045 28.432 4.236 4.417 35.950 3.916 0.794 2.326 0.800 2.319
Office1 43.651 31.719 62.805 42.110 11.592 4.145 60.846 3.468 0.778 3.625 0.740 3.426
Office2 24.612 190.52 44.780 128.35 0.526 4.870 29.353 12.435 0.442 3.156 0.419 2.980
Room0 44.478 173.42 32.990 229.53 0.497 2.343 38.550 13.274 0.889 8.321 0.893 8.308
Room1 38.663 243.30 44.063 109.78 0.306 2.411 34.774 5.936 0.332 1.760 0.318 1.695
DTU scan24 23.123 10.736 4.830 30.796 1.501 3.564 29.952 1.880 0.147 0.383 0.152 0.388
scan37 56.389 95.731 10.628 8.832 5.295 18.280 47.159 15.749 0.461 1.446 0.441 1.458
scan40 50.555 70.607 6.443 15.590 1.289 5.681 59.713 7.021 0.626 1.418 0.628 1.438
scan55 30.339 69.516 4.707 15.411 1.325 11.293 49.758 4.359 0.203 1.209 0.216 1.221
scan63 39.299 79.984 3.520 30.415 4.760 17.137 37.742 17.378 0.530 2.206 0.516 2.203
scan65 11.181 94.896 6.210 39.433 2.124 16.450 25.381 12.917 0.230 1.241 0.276 1.125
scan69 54.851 72.511 2.941 13.699 1.896 6.562 41.689 7.427 0.128 1.853 0.092 1.826
scan83 42.331 86.062 1.877 12.644 6.083 36.981 77.169 26.918 0.325 2.530 0.487 2.181
scan97 29.318 67.483 16.416 30.376 2.506 12.922 56.655 13.913 0.308 1.073 0.313 1.069
scan105 54.325 67.535 3.267 26.114 3.190 26.857 34.217 8.835 0.219 2.054 0.210 2.108
scan106 54.349 72.003 5.200 8.806 0.460 10.196 41.294 4.361 0.359 1.196 0.366 1.215
scan110 40.470 61.468 10.383 13.227 1.909 7.209 42.506 8.192 0.296 0.810 0.301 0.761
scan114 8.837 61.963 5.450 38.759 1.854 10.781 26.354 10.068 0.357 1.123 0.363 1.086
scan118 43.873 81.839 2.640 3.735 2.032 15.600 35.484 10.656 0.081 0.576 0.089 0.502
scan122 34.831 71.341 4.121 23.594 2.499 13.661 39.012 8.106 0.244 0.366 0.257 0.258
Table S5: Detailed pose estimation experiments on Replica and DTU datasets.